The consumption rate of a certain item in a company is 100 units/day. The ordering cost per order is Rs.100 and the average daily holding cost per unit is 2 paise. Determine the economic order quantity for placing an order. Also, find the total cost per unit time (TCU) and the cycle time.
1
Expert's answer
2019-01-26T03:47:07-0500

D = 100 units/day, S = Rs.100, H = 0.02.

The economic order quantity for placing an order is:

EOQ = (2*D*S/H)^0.5 = (2*100*100/0.02)^0.5 = 1000.

The total cost per unit time is:

TCU = SD/Q + HQ/2 = 100*100/1000 + 0.02*1000/2 = 20.

The cycle time is:

t = Q/D = 1000/100 = 10.